# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

### v1

@@ Appendix E Landscape Similarity & Datasets Characteristics
While in general, we find that HP loss landscapes studied in this paper share various common characteristics, the detailed topography can vary with dataset that the model is trained on. How characteristics of datasets could affect the landscape is an interesting problem to explore in more detail. In addition, we also hypothesis that the generalization gap of a model can also depend on the dataset, in addition to its HP setting. To investigate these, we conduct additional experiments on the $57$ tabular datasets to analyze the relationship between landscape similarity and dataset characteristics.
@@
We first find in the left plot of Figure 15 that the similarity between test and training loss landscapes is positively correlated across models. It implies that on certain datasets, all the $4$ models are more prone to overfit, whereas the opposite could be true for other datasets. This then verifies our hypothesis that overfitting not only depends on the model HPs, but also the dataset itself. To further explore which properties of the dataset could potentially contribute to this, we investigated the correlation between train-test landscape similarity and the number of instances & features of each dataset, and their product. The results (also in Figure 15 (left)) indicate that the degree of overfitting is correlated with all $3$ dataset size measures. In particular, for most models, it would be more likely to encounter overfitting on larger datasets.
@@
We then proceed to investigate whether correlations between HP loss landscapes induced on different datasets are related to the relative size of the datasets. From Figure 15 (right), it is clear to see that pairwise landscape similarities are obviously correlated across models, implying that all the models are likely to induce very different (or the opposite) landscapes on certain pairs of datasets. We could also see that these pairwise similarities are again correlated with the differences in dataset sizes. In particular, for datasets that have very different sizes, the resulting HP loss landscapes also tend to be different from each other.
@@
However, we note that the correlations reported in both scenarios are somewhat weak, and the reasons could be two folds. First, the choice of datasets is only a partial factor contributing to landscape similarity, many other factors like model HPs, training settings can also important roles here. Second, the dataset meta-features we used here are rather naive, and there could be more comprehensive features for characterizing dataset properties, e.g., Feurer et al. (2015b) had used a collection of $46$ features to assess dataset similarity. However, this is beyond the scope of this work and we leave it to future works.
Appendix F Details of the Experimental Setup
F.1 Search Spaces
@@
In this subsection, we elaborate the principles that we follow in designing the search spaces for each model. We also provide the detailed hyperparameter grid space for them in Table 4 (XGBoost), Table 5 (DT), Table 6 (RF), Table 7 (RF), Table 8 (CNN) and Table 9 (FCNet). The top-level principle we follow is to include commonly used HPs and exclude unimportant ones, while keeping a good balance between search space coverage and computational cost. We first determine the list of HPs to be considered, and to this end:
@@
  • We surveyed commonly used HPs in the practice (e.g., van Rijn & Hutter identified important HPs of several models using large-scale meta-data from OpenML) and HPO literature (e.g., the search spaces used in HPOBench (Eggensperger et al., 2021)). For CNN, we also refer to the design of NAS search spaces (Chitty-Venkata et al., 2023).
@@
width_linear $[256,512,1024]$ $3$
width_conv $[32,64,128]$ $3$
#conv_block $[2,4,6]$ $3$
Table 9: FCNet HP Grid Space ($62,208$ Configurations)
 Hyperparameter Grid Values Count
initial_LR $\{5\mathrm{e}^{-4},1\mathrm{e}^{-3},5\mathrm{e}^{-3},1\mathrm{e}^{-2},1\mathrm{e}^{-2},.1\mathrm{e}^{-1}\}$ $6$
batch_size $\{8,16,32,64\}$ $4$
LR_schedule {cosine, fix} $2$
activation_layer_1 {relu, tanh} $2$
activation_layer_2 {relu, tanh} $2$
layer_1_Size $\{16,32,64,128,256,512\}$ $6$
layer_2_Size $\{16,32,64,128,256,512\}$ $6$
dropout_Layer_1 $\{0.0,0.3,0.6\}$ $3$
dropout_Layer_2 $\{0.0,0.3,0.6\}$ $3$
F.2 Datasets

### v2

@@ Appendix F Landscape Similarity & Datasets Characteristics
While in general, we find that HP loss landscapes studied in this paper share various common characteristics, the detailed topography can vary with dataset that the model is trained on. How characteristics of datasets could affect the landscape is an interesting problem to explore in more detail. In addition, we also hypothesize that the generalization gap of a model can also depend on the dataset, in addition to its HP setting. To investigate these, we conduct additional experiments on the $57$ tabular datasets to analyze the relationship between landscape similarity and dataset characteristics.
@@
We first find in the left plot of Fig. 15 that the similarity between test and training loss landscapes is positively correlated across models. It implies that on certain datasets, all the $4$ models are more prone to overfit, whereas the opposite could be true for other datasets. This then verifies our hypothesis that overfitting not only depends on the model HPs, but also the dataset itself. To further explore which properties of the dataset could potentially contribute to this, we investigated the correlation between train-test landscape similarity and the number of instances & features of each dataset, and their product. The results (also in Fig. 15 (left)) indicate that the degree of overfitting is correlated with all $3$ dataset size measures. In particular, for most models, it would be more likely to encounter overfitting on larger datasets.
@@
We then proceed to investigate whether correlations between HP loss landscapes induced on different datasets are related to the relative size of the datasets. From Fig. 15 (right), it is clear to see that pairwise landscape similarities are obviously correlated across models, implying that all the models are likely to induce very different (or the opposite) landscapes on certain pairs of datasets. We could also see that these pairwise similarities are again correlated with the differences in dataset sizes. In particular, for datasets that have very different sizes, the resulting HP loss landscapes also tend to be different from each other.
@@
However, we note that the correlations reported in both scenarios are somewhat weak, and the reasons could be two folds. First, the choice of datasets is only a partial factor contributing to landscape similarity, many other factors like model HPs, training settings can also important roles here. Second, the dataset meta-features we used here are rather naive, and there could be more comprehensive features for characterizing dataset properties, e.g., [53] had used a collection of $46$ features to assess dataset similarity. However, this is beyond the scope of this work and we leave it to future works.
Appendix G Details of the Experimental Setup
G.1 Search Spaces
@@
In this subsection, we elaborate the principles that we follow in designing the search spaces for each model. We also provide the detailed hyperparameter grid space for them in Table 2 (XGBoost), Table 3 (DT), Table 4 (RF), Table 5 (RF), Table 6 (SVM), and Table 7 (CNN). The top-level principle we follow is to include commonly used HPs and exclude unimportant ones, while keeping a good balance between search space coverage and computational cost. We first determine the list of HPs to be considered, and to this end:
@@
  • We surveyed commonly used HPs in the practice (e.g., [87] identified important HPs of several models using large-scale meta-data from OpenML) and HPO literature (e.g., the search spaces used in HPOBench [88]). For CNN, we also refer to the design of NAS search spaces [89].
@@
Table 6: SVM HP Grid Space ($2340$ Configurations)
Hyperparameter Grid Values Count
gamma $[1\mathrm{e}^{-4},10]$ $45$
c $[1\mathrm{e}^{-3},\dots,1\mathrm{e}^{3}]$ $54$
@@
width_linear [$256,512,1024$] $3$
width_conv [$32,64,128$] $3$
#conv_block [$2,4,6$] $3$
G.2 Datasets